# Bangle functions are the generic basis for cluster algebras from punctured surfaces with boundary

### v2

@@ 3.2. KRS-monoids of $\tau^{-}$-reduced components
:$\oplus:\{(X,Y)\in\operatorname{DecIrr}^{\tau^{-}}(A)\times\operatorname{DecIrr}^{\tau^{-}}(A)\mid e(x,y)=0\}\rightarrow\operatorname{DecIrr}^{\tau^{-}}(A)$
@@
    In the form it is stated, Theorem 3.6 appeared (up to duality) in [24]. Its assertions are restatements of previous results by various authors: Part (a) is a well-known combination of [7, Theorem 1.2] and [6, Theorems 1.3 and 1.5]. Part (b) is a theorem of Plamondon [47]. Part (c) is [26, Corollary 1.7] and [23, Theorem 3.2].
@@
    More precisely, Theorem 3.6 is just the dual version of [24, Thm. 2.5]. There, the result was stated in terms of generically $\tau$-reduced components with the function $e_{A}^{\operatorname{proj}}$ which is defined in terms of the projective symmetrized E-invariant
@@
    Duality over the ground field $𝕜\Bbbk$ induces a natural isomorphism of framed partial KRS-monoids
@@
In view of [10, Proposition 10.4 and Remark 10.8], we have
@@
Here $\mathbf{g}_{A}^{\mathrm{DWZ}}(M)$ denotes the $g$-vector of a QP-representation in the sense of [10, (1.13)]. This can be easily extended to decorated QP-representations.
@@
where $B(Q)=(b_{ij})$ is the matrix defined in (1.2). Note that, this is just another way of writing the transformation rule for g-vectors from [19, Conj. 7.12], which was proved in [10, Thm. 1.7]. Compare also with Equation (5.6).
@@
[24, Proposition 2.8] For each $Z\in\operatorname{DecIrr}^{\tau^{-}}(A)$ there exists a dense open subset $U_{Z}\subset Z$, a unique irreducible component $\tilde{\mu}_{k}(Z)\in\operatorname{DecIrr}^{\tau^{-}}(A^{\prime})$ and a regular map $\nu_{Z}:U_{Z}\rightarrow\tilde{\mu}_{k}(Z)$ with the following properties:
@@
of $(n+1)\times n$ matrices, where $\mu_{k}$ is Fomin-Zelevinsky’s matrix mutation. Thus, Proposition 3.8 is a generalization of Nakanishi-Zelevinsky’s result [45, Equation (4.1)].
@@ 3.3. KRS-monoids of laminations
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a possibly-punctured surface with marked points. In [24, Section 4], we considered the set $\mathcal{LC}^{*}(\mathbf{\Sigma})$ of homotopy classes of marked curves and loops on $\mathbf{\Sigma}$ that have no kinks. Inspired by [50], we introduce a symmetric marked intersection number
@@
Given a marked curve (or loop) $(\gamma,c)\in\mathcal{LC}^{*}(\mathbf{\Sigma})$, let $(\gamma,c)^{-1}$ denote the inversely oriented curve $\gamma$, with accordingly swapped decoration $c$. The involution $(\gamma,c)\mapsto(\gamma,c)^{-1}$ induces an equivalence relation $\simeq$ on $\mathcal{LC}^{*}(\mathbf{\Sigma})$. For us, as in [50], a simple marked curve is a marked curve $(\gamma,c)\in\mathcal{LC}^{*}(\mathbf{\Sigma})$ which has self-intersection number 00. We let $\mathcal{LC}^{*}_{\tau}(\mathbf{\Sigma})$ denote the set of all simple marked curves, and define the tame partial KRS-monoid of laminations
@@
See Example 3.4 and [24, Section 4.3] for more details. The set $\mathcal{LC}^{*}_{\tau}(\mathbf{\Sigma})/_{\simeq}$ can be identified with the set of laminations $\mathcal{C}^{\circ}(\mathbf{\Sigma})$ considered in [44] (two marked curves having intersection number equal to 00 corresponds to them being $\mathcal{C}^{\circ}$-compatible in Musiker-Schiffler-Williams’ nomenclature). This is also compatible with the treatment of laminations in [16].
@@
Following [15], represent $T$ through an ideal triangulation $T^{\circ}$ defined as follows. For each $p\in\mathbb{P}$ such that $\delta_{T}(p)=0$, let $i_{p},j_{p},$ be the two tagged arcs in $T$ that are incident to $p$, with $i_{p}$ tagged plain at $p$, and $j_{p}$ tagged notched at $p$. Then $T^{\circ}:=\{\gamma^{\circ}\ |\ \gamma\in T\}$ is obtained from $T$ by setting $\gamma^{\circ}:=\gamma$ for every $\gamma\in T$ both of whose ends are tagged plain, and by replacing each $j_{p}$ with a loop $j_{p}^{\circ}$ closely surrounding $i_{p}^{\circ}$ for $i$ and $j$ as above. Thus, the corresponding ordinary arcs $i_{p}^{\circ},j_{p}^{\circ}\in T^{\circ}$ form a self folded triangle that has $i_{p}^{\circ}$ as the folded side, and $j_{p}^{\circ}$ as the loop closely surrounding $i_{p}^{\circ}$. See [15, Sections 9.1 and 9.2].
@@
Suppose that the ordinary arc $\gamma^{\circ}\in T^{\circ}$ is not the folded side of a self-folded triangle of $T^{\circ}$. Then $\gamma^{\circ}$ is contained in exactly two ideal triangles of $T^{\circ}$, and the union ¯$\overline{\lozenge}_{\gamma^{\circ}}$ of these two triangles is either a quadrilateral (if $\gamma^{\circ}$ does not enclose a self-folded triangle) or a digon (if $\gamma^{\circ}$ encloses a self-folded triangle). In any of these two situations, the complement in ¯$\overline{\lozenge}_{\gamma^{\circ}}$ of the union of the arcs belonging to $T^{\circ}\setminus\{\gamma^{\circ}\}$ can be thought to be an open quadrilateral $\lozenge_{\gamma^{\circ}}$ in which $\gamma^{\circ}$ sits as a diagonal. The shear coordinate $\operatorname{Sh}_{T}(\lambda)_{\gamma}$ is defined to be the number of segments of $\lozenge_{\gamma^{\circ}}\cap L$ that form the shape of a letter $Z$ when crossing $\gamma^{\circ}$ minus the number of segments of $\lozenge_{\gamma^{\circ}}\cap L$ that form the shape of a letter $S$ when crossing $\gamma^{\circ}$.
@@
where $\operatorname{Sh}_{T}(\lambda)$ is the vector of dual shear coordinates we have defined above, $\overline{T}$ and $\overline{\lambda}$ are the images of $T$ and $\lambda$ in the surface obtained as the mirror image of $(\Sigma,\mathbb{M},\mathbb{P})$, and $\operatorname{Sh}_{\overline{T}}^{\operatorname{FST}}(\overline{\lambda})$ is the vector of shear coordinates of Fomin–Thurston, [16, Definition 13.1].
With these definitions in place we can restate [16, Theorem 13.5] for dual shear coordinates as follows.
@@
Generalizing [23, Theorem 10.13 and Proposition 10.14] from the unpunctured to the punctured case, the main result of [24] states with our conventions
@@
[24, Theorem 1.1.] Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with non-empty boundary. For each tagged triangulation $T$ of $\mathbf{\Sigma}$ there is a unique isomorphism of partial KRS-monoids
@@
[…]
where $A(T)$ is the dual Jacobian algebra of the quiver with potential associated to $T$ in [5, Definition 4.1], [36, Definitions 3.1 and 3.2], and $\mathbf{g}_{A(T)}$ is the generic injective g-vector.
@@
    the statement of the theorem is shown to hold for tagged triangulations of signature zero, exploiting heavily the fact that for those triangulations the Jacobian algebra is skewed-gentle, and that in this situation a combinatorial description of the generically $\tau^{-}$-reduced components is now available [20];
@@ 4. Triangulations adapted to closed curves
Let $\alpha$ be an essential loop on $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$, i.e., a non-contractible simple closed curve that is furthermore not contractible to a puncture, and let $T=\{\tau_{1},\ldots,\tau_{n}\}$ be a tagged triangulation of $\mathbf{\Sigma}$. The band graph associated to $\alpha$ with respect to $T$, cf. [44, Definition 3.4 and §8.3], will be denoted $G(T,\alpha)$. The polynomial $F_{\zeta}^{T}$ defined in [44, Definition 3.14] (see also [44, §8.3]), with $\zeta:=\alpha$, will be called snake $F$-polynomial and denoted $F_{G(T,\alpha)}$. Similarly, the integer vector from [44, Definition 6.1] will be called the snake $g$-vector and denoted $\mathbf{g}_{G(T,\alpha)}$. Notice that
@@
where $\mathcal{P}(G(T,\alpha))$ is the collection of all good matchings of the band graph $G(T,\alpha)$. Furthermore, by [44, Proposition 6.2], we have
@@
where $P_{-}$ is the minimal matching of the band graph $G(T,\alpha)$, cf. [44, Definition 3.7], and $\deg:\mathbb{Z}[x_{1}^{\pm 1},\ldots,x_{n}^{\pm 1},y_{1},\ldots,y_{n}]\rightarrow\mathbb{Z}^{n}$ is the grading defined by
@@
as in [19, Proposition 6.1], $\mathbf{b}(T)_{j}$ being the $j^{\operatorname{th}}$ column of $B(T)$, see Definition 2.9.
@@
In both cases, it is clear that $\Lambda_{I}:=\mathcal{P}(Q(T_{\alpha})|_{I},S(T_{\alpha})|_{I})$ is a gentle algebra. That $M(T_{\alpha},\alpha,\lambda,1)$ is a module also over $A(T_{\alpha})$ follows immediately from [9, Proposition 8.9] or by direct inspection.
It is easy to check that $A(\overline{T_{\alpha}})=A(T_{\alpha})^{\operatorname{op}}$ and $M(\overline{T_{\alpha}},\overline{\alpha},\lambda,1)\cong D(M(T_{\alpha},\alpha,\lambda,1))$ as representations of $A(\overline{T_{\alpha}})$. Therefore, by (3.2), [34, proof of Theorem 10.0.5] and (3.4), we have
@@
The equality $\operatorname{Sh}_{T_{\alpha}}(\alpha)=\mathbf{g}_{G(T_{\alpha},\alpha)}$ follows from [23, Proposition 10.14].
@@
${\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$${\ldots}{\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$$I$$I$$I$$I$$\lambda^{-1}$
Now, we can use Krause’s formalism from [33]. With the notation from loc. cit. p. 191 we denote by $\mathcal{A}(b,b)$ the set of isomorphism classes of admissable (sic) triples for the pair of bands $(b,b)$. Since the band $b$ has a unique source, which is mapped to $t_{i_{1}}$ and a unique sink, which is mapped to $t_{i_{n}}\neq t_{i_{1}}$ we have $\mathcal{A}(b,b)=\{[(b,\operatorname{id}_{b},\operatorname{id}_{b})]\}$. Thus, $\operatorname{End}_{\Lambda_{I}}(M)\cong\mathbb{C}$ follows directly from the Theorem on [33, p. 191]. ∎
@@
The equality $F_{M(T_{\alpha},\alpha,\lambda,1)}=F_{G(T_{\alpha},\alpha)}$ follows from [23, Lemma 11.4 and Remark 11.7] and [29, Theorem 1.2] as in the proof of [23, Proposition 11.8]. Note however, that due to our convention for $A(T_{\alpha})$ resp. for $\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T_{\alpha})\mid_{I},S^{\prime}(T_{\alpha})_{I})$, we can avoid the use of dual CC-functions. This finishes the proof of Proposition 4.1. ∎
@@
    is a generically $\tau^{-}$-reduced indecomposable irreducible component of the representation space $\operatorname{rep}(A(T_{\alpha}),\mathbf{d})$, where $\mathbf{d}:=\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))$;
@@
By [21, Theorem 7.1] (see also [35, Theorem 3.5]), the algebra $A(T_{\alpha})$ is tame. Furthermore, for $\lambda_{1}\neq\lambda_{2}$, the band modules $M(T_{\alpha},\alpha,\lambda_{1},1)$ and $M(T_{\alpha},\alpha,\lambda_{2},1)$ are non-isomorphic, and by Proposition 4.1, for every point $M$ in the dense open subset $\bigcup_{\lambda\in\mathbb{C}\setminus\{0\}}\operatorname{GL}_{\mathbf{d}}(\mathbb{C})\cdot M(T_{\alpha},\alpha,\lambda,1)$ we have $E_{A(T_{\alpha})}(M)=1$. Hence $Z_{T_{\alpha},\alpha}$ is a generically $\tau^{-}$-reduced indecomposable irreducible component by [23, Lemma 3.1 and Theorem 3.2] (see also [3, Section 2.2]).

### v3

@@ 3.2. KRS-monoids of $\tau^{-}$-regular components
$\oplus:\{(X,Y)\in\operatorname{DecIrr}^{\tau^{-}}(A)\times\operatorname{DecIrr}^{\tau^{-}}(A)\mid e(x,y)=0\}\rightarrow\operatorname{DecIrr}^{\tau^{-}}(A)$
@@
    In the form it is stated, Theorem 3.6 appeared (up to duality) in [25]. Its assertions are restatements of previous results by various authors: Part (a) is a well-known combination of [8, Theorem 1.2] and [7, Theorems 1.3 and 1.5]. Part (b) is a theorem of Plamondon [48]. Part (c) is [27, Corollary 1.7] and [24, Theorem 3.2].
@@
    More precisely, Theorem 3.6 is just the dual version of [25, Thm. 2.5]. There, the result was stated in terms of generically $\tau$-regular components with the function $e_{A}^{\operatorname{proj}}$ which is defined in terms of the projective symmetrized E-invariant
@@
[…]
@@
In view of [11, Proposition 10.4 and Remark 10.8], we have
@@
Here $\mathbf{g}_{A}^{\mathrm{DWZ}}(M)$ denotes the $g$-vector of a QP-representation in the sense of [11, (1.13)]. This can be easily extended to decorated QP-representations.
@@
where $B(Q)=(b_{ij})$ is the matrix defined in (1.2). Note that, this is just another way of writing the transformation rule for g-vectors from [20, Conj. 7.12], which was proved in [11, Thm. 1.7]. Compare also with Equation (5.6).
@@
[25, Proposition 2.8] For each $Z\in\operatorname{DecIrr}^{\tau^{-}}(A)$ there exists a dense open subset $U_{Z}\subset Z$, a unique irreducible component $\tilde{\mu}_{k}(Z)\in\operatorname{DecIrr}^{\tau^{-}}(A^{\prime})$ and a regular map $\nu_{Z}:U_{Z}\rightarrow\tilde{\mu}_{k}(Z)$ with the following properties:
@@
of $(n+1)\times n$ matrices, where $\mu_{k}$ is Fomin-Zelevinsky’s matrix mutation. Thus, Proposition 3.8 is a generalization of Nakanishi-Zelevinsky’s result [46, Equation (4.1)].
@@ 3.3. KRS-monoids of laminations
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a possibly-punctured surface with marked points. In [25, Section 4], we considered the set $\mathcal{LC}^{*}(\mathbf{\Sigma})$ of homotopy classes of marked curves and loops on $\mathbf{\Sigma}$ that have no kinks. Inspired by [51], we introduce a symmetric marked intersection number
@@
Given a marked curve (or loop) $(\gamma,c)\in\mathcal{LC}^{*}(\mathbf{\Sigma})$, let $(\gamma,c)^{-1}$ denote the inversely oriented curve $\gamma$, with accordingly swapped decoration $c$. The involution $(\gamma,c)\mapsto(\gamma,c)^{-1}$ induces an equivalence relation $\simeq$ on $\mathcal{LC}^{*}(\mathbf{\Sigma})$. For us, as in [51], a simple marked curve is a marked curve $(\gamma,c)\in\mathcal{LC}^{*}(\mathbf{\Sigma})$ which has self-intersection number 0. We let $\mathcal{LC}^{*}_{\tau}(\mathbf{\Sigma})$ denote the set of all simple marked curves, and define the tame partial KRS-monoid of laminations
@@
See Example 3.4 and [25, Section 4.3] for more details. The set $\mathcal{LC}^{*}_{\tau}(\mathbf{\Sigma})/_{\simeq}$ can be identified with the set of laminations $\mathcal{C}^{\circ}(\mathbf{\Sigma})$ considered in [45] (two marked curves having intersection number equal to 0 corresponds to them being $\mathcal{C}^{\circ}$-compatible in Musiker-Schiffler-Williams’ nomenclature). This is also compatible with the treatment of laminations in [17].
@@
Following [16], represent $T$ through an ideal triangulation $T^{\circ}$ defined as follows. For each $p\in\mathbb{P}$ such that $\delta_{T}(p)=0$, let $i_{p},j_{p},$ be the two tagged arcs in $T$ that are incident to $p$, with $i_{p}$ tagged plain at $p$, and $j_{p}$ tagged notched at $p$. Then $T^{\circ}:=\{\gamma^{\circ}\ |\ \gamma\in T\}$ is obtained from $T$ by setting $\gamma^{\circ}:=\gamma$ for every $\gamma\in T$ both of whose ends are tagged plain, and by replacing each $j_{p}$ with a loop $j_{p}^{\circ}$ closely surrounding $i_{p}^{\circ}$ for $i$ and $j$ as above. Thus, the corresponding ordinary arcs $i_{p}^{\circ},j_{p}^{\circ}\in T^{\circ}$ form a self folded triangle that has $i_{p}^{\circ}$ as the folded side, and $j_{p}^{\circ}$ as the loop closely surrounding $i_{p}^{\circ}$. See [16, Sections 9.1 and 9.2].
@@
Suppose that the ordinary arc $\gamma^{\circ}\in T^{\circ}$ is not the folded side of a self-folded triangle of $T^{\circ}$. Then $\gamma^{\circ}$ is contained in exactly two ideal triangles of $T^{\circ}$, and the union $\overline{\lozenge}_{\gamma^{\circ}}$ of these two triangles is either a quadrilateral (if $\gamma^{\circ}$ does not enclose a self-folded triangle) or a digon (if $\gamma^{\circ}$ encloses a self-folded triangle). In any of these two situations, the complement in $\overline{\lozenge}_{\gamma^{\circ}}$ of the union of the arcs belonging to $T^{\circ}\setminus\{\gamma^{\circ}\}$ can be thought to be an open quadrilateral $\lozenge_{\gamma^{\circ}}$ in which $\gamma^{\circ}$ sits as a diagonal. The shear coordinate $\operatorname{Sh}_{T}(\lambda)_{\gamma}$ is defined to be the number of segments of $\lozenge_{\gamma^{\circ}}\cap L$ that form the shape of a letter $Z$ when crossing $\gamma^{\circ}$ minus the number of segments of $\lozenge_{\gamma^{\circ}}\cap L$ that form the shape of a letter $S$ when crossing $\gamma^{\circ}$.
@@
where $\operatorname{Sh}_{T}(\lambda)$ is the vector of dual shear coordinates we have defined above, $\overline{T}$ and $\overline{\lambda}$ are the images of $T$ and $\lambda$ in the surface obtained as the mirror image of $(\Sigma,\mathbb{M},\mathbb{P})$, and $\operatorname{Sh}_{\overline{T}}^{\operatorname{FST}}(\overline{\lambda})$ is the vector of shear coordinates of Fomin–Thurston, [17, Definition 13.1].
With these definitions in place we can restate [17, Theorem 13.5] for dual shear coordinates as follows.
@@
Generalizing [24, Theorem 10.13 and Proposition 10.14] from the unpunctured to the punctured case, the main result of [25] states with our conventions
@@
[25, Theorem 1.1.] Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with non-empty boundary. For each tagged triangulation $T$ of $\mathbf{\Sigma}$ there is a unique isomorphism of partial KRS-monoids
@@
${\operatorname{Lam}(\mathbf{\Sigma})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{T}}$$\scriptstyle{\operatorname{Sh}_{T}}$${\operatorname{DecIrr}^{\tau^{-}}(A(T))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{g}_{A(T)}}$${\mathbb{Z}^{T}}$
where $A(T)$ is the dual Jacobian algebra of the quiver with potential associated to $T$ in [6, Definition 4.1], [37, Definitions 3.1 and 3.2], and $\mathbf{g}_{A(T)}$ is the generic injective g-vector.
@@
    the statement of the theorem is shown to hold for tagged triangulations of signature zero, exploiting heavily the fact that for those triangulations the Jacobian algebra is skewed-gentle, and that in this situation a combinatorial description of the generically $\tau^{-}$-regular components is now available [21];
@@ 4. Triangulations adapted to closed curves
Let $\alpha$ be an essential loop on $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$, i.e., a non-contractible simple closed curve that is furthermore not contractible to a puncture, and let $T=\{\tau_{1},\ldots,\tau_{n}\}$ be a tagged triangulation of $\mathbf{\Sigma}$. The band graph associated to $\alpha$ with respect to $T$, cf. [45, Definition 3.4 and §8.3], will be denoted $G(T,\alpha)$. The polynomial $F_{\zeta}^{T}$ defined in [45, Definition 3.14] (see also [45, §8.3]), with $\zeta:=\alpha$, will be called snake $F$-polynomial and denoted $F_{G(T,\alpha)}$. Similarly, the integer vector from [45, Definition 6.1] will be called the snake $g$-vector and denoted $\mathbf{g}_{G(T,\alpha)}$. Notice that
@@
where $\mathcal{P}(G(T,\alpha))$ is the collection of all good matchings of the band graph $G(T,\alpha)$. Furthermore, by [45, Proposition 6.2], we have
@@
where $P_{-}$ is the minimal matching of the band graph $G(T,\alpha)$, cf. [45, Definition 3.7], and $\deg:\mathbb{Z}[x_{1}^{\pm 1},\ldots,x_{n}^{\pm 1},y_{1},\ldots,y_{n}]\rightarrow\mathbb{Z}^{n}$ is the grading defined by
@@
as in [20, Proposition 6.1], $\mathbf{b}(T)_{j}$ being the $j^{\operatorname{th}}$ column of $B(T)$, see Definition 2.9.
@@
In both cases, it is clear that $\Lambda_{I}:=\mathcal{P}(Q(T_{\alpha})|_{I},S(T_{\alpha})|_{I})$ is a gentle algebra. That $M(T_{\alpha},\alpha,\lambda,1)$ is a module also over $A(T_{\alpha})$ follows immediately from [10, Proposition 8.9] or by direct inspection.
It is easy to check that $A(\overline{T_{\alpha}})=A(T_{\alpha})^{\operatorname{op}}$ and $M(\overline{T_{\alpha}},\overline{\alpha},\lambda,1)\cong D(M(T_{\alpha},\alpha,\lambda,1))$ as representations of $A(\overline{T_{\alpha}})$. Therefore, by (3.2), [35, proof of Theorem 10.0.5] and (3.4), we have
@@
The equality $\operatorname{Sh}_{T_{\alpha}}(\alpha)=\mathbf{g}_{G(T_{\alpha},\alpha)}$ follows from [24, Proposition 10.14].
@@
(4.1) $t_{i_{n}}$$t_{i_{n-1}}$$t_{i_{n-2}}$$\ldots$$t_{i_{3}}$$t_{i_{2}}$$t_{i_{1}}$$\alpha_{n-1}$$\alpha_{n-2}$$\alpha_{2}$$\alpha_{1}$$\alpha_{n}$
@@
Now, we can use Krause’s formalism from [34]. With the notation from loc. cit. p. 191 we denote by $\mathcal{A}(b,b)$ the set of isomorphism classes of admissable (sic) triples for the pair of bands $(b,b)$. Since the band $b$ has a unique source, which is mapped to $t_{i_{1}}$ and a unique sink, which is mapped to $t_{i_{n}}\neq t_{i_{1}}$ we have $\mathcal{A}(b,b)=\{[(b,\operatorname{id}_{b},\operatorname{id}_{b})]\}$. Thus, $\operatorname{End}_{\Lambda_{I}}(M)\cong\mathbb{C}$ follows directly from the Theorem on [34, p. 191]. ∎
@@
The equality $F_{M(T_{\alpha},\alpha,\lambda,1)}=F_{G(T_{\alpha},\alpha)}$ follows from [24, Lemma 11.4 and Remark 11.7] and [30, Theorem 1.2] as in the proof of [24, Proposition 11.8]. Note however, that due to our convention for $A(T_{\alpha})$ resp. for $\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T_{\alpha})\mid_{I},S^{\prime}(T_{\alpha})_{I})$, we can avoid the use of dual $C$-functions. This finishes the proof of Proposition 4.1. ∎
@@
    is a generically $\tau^{-}$-regular indecomposable irreducible component of the representation space $\operatorname{rep}(A(T_{\alpha}),\mathbf{d})$, where $\mathbf{d}:=\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))$;
@@
By [22, Theorem 7.1] (see also [36, Theorem 3.5]), the algebra $A(T_{\alpha})$ is tame. Furthermore, for $\lambda_{1}\neq\lambda_{2}$, the band modules $M(T_{\alpha},\alpha,\lambda_{1},1)$ and $M(T_{\alpha},\alpha,\lambda_{2},1)$ are non-isomorphic, and by Proposition 4.1, for every point $M$ in the dense open subset $\bigcup_{\lambda\in\mathbb{C}\setminus\{0\}}\operatorname{GL}_{\mathbf{d}}(\mathbb{C})\cdot M(T_{\alpha},\alpha,\lambda,1)$ we have $E_{A(T_{\alpha})}(M)=1$. Hence $Z_{T_{\alpha},\alpha}$ is a generically $\tau^{-}$-regular indecomposable irreducible component by [24, Lemma 3.1 and Theorem 3.2] (see also [4, Section 2.2]).